OpenStudy (prince1342):

Simplify completely quantity 7 x squared plus 11 x minus 6 all over quantity 7 x squared minus 10 x plus 3.

OpenStudy (prince1342):

quantity x plus 2 over quantity 7 x plus 3 quantity x plus 2 over quantity x minus 1 quantity 7 x minus 6 over quantity x minus 1 quantity 7 x minus 6 over quantity 7 x plus 3

OpenStudy (welshfella):

7x^2 + 11x - 6 -------------- 7x^2 - 10x + 3

OpenStudy (prince1342):

Yep! Thats what it looks like

OpenStudy (welshfella):

can you factor the 2 functions?

OpenStudy (welshfella):

7x^2 + 11x - 6 = (7x + a)(x + b) we need to figure out what a and b might be

OpenStudy (prince1342):

i dont know how to factor them

OpenStudy (welshfella):

a times b must equal -6 and a + 7b must equal 11

OpenStudy (welshfella):

lets try the ac method 7 * 6 = -42 so we have to look at the factors of -42

OpenStudy (prince1342):

how did you get negative by multiplying 2 positives?

OpenStudy (welshfella):

- and keep in mind that miiddle term is + 11x

OpenStudy (welshfella):

7*-6 sorry

OpenStudy (prince1342):

oh ok lol

OpenStudy (prince1342):

Oh! So ac represent 7 and -6

OpenStudy (welshfella):

14* -3 = -42 looks promisiing because tis will give us 14 - 3 = 11

OpenStudy (welshfella):

yea from ax^2 + bx + c

OpenStudy (welshfella):

so lets try 7x^2 + 14 - 3x - 6

OpenStudy (welshfella):

factior this by grouping ( that is the first 2 and last 2 terms)

OpenStudy (prince1342):

how did you get 14?

OpenStudy (welshfella):

7x^2 + 14x - 3x - 6 = 7x(x + 2) - 3(x + 2) do you follow that OK?

OpenStudy (welshfella):

- i should have written 14x

OpenStudy (welshfella):

typo

OpenStudy (prince1342):

ok i follow

OpenStudy (welshfella):

ok now since x+ 2 is common to the 2 parts we can write (x + 2)(7x - 3) and theses are the factors of the top part of the fraction 7x^2 - 10x + 3 can be factored in a similar fashion. note , here ac = + 21

OpenStudy (welshfella):

now to get -10x you need -7 and -3 so can you continue?

OpenStudy (welshfella):

its -7 and -3 because their product is 21 and theire sum is -10

OpenStudy (prince1342):

so 3 over 7

OpenStudy (welshfella):

no - next its 7x^2 - 7x - 3x + 3 - now you need to factor by grouping as we did before

OpenStudy (welshfella):

fist factor 7x^2 - 7x ( take out common factor 7x)

OpenStudy (welshfella):

- look back at the previous post where we were at this stage

OpenStudy (prince1342):

Im lost :/

OpenStudy (welshfella):

divide 7x^2 by 7x then divide -7x by 7x and youll get what goes in the parenthesis

OpenStudy (welshfella):

7x^2 - 7x = 7x(x - 1)

OpenStudy (welshfella):

so we have 7x(x - 1) - 3x + 3 - now factor -3x - 3 ( take out 3)

OpenStudy (welshfella):

* take out -3

OpenStudy (welshfella):

still confused?

OpenStudy (prince1342):

yeah

OpenStudy (prince1342):

So many numbers

OpenStudy (welshfella):

-3x - 3 = -3(x + 1) - understand this?

OpenStudy (prince1342):

yes

OpenStudy (welshfella):

-3x / -3 = x and -3/-3 = + 1 so x + 1 goes into the brackets. giving -3(x + 1)

OpenStudy (welshfella):

ok so we have 7x(x + 1) - 3(x + 1) - now (x +1) is common to both parts so it simplifies to (7x - 3)(x + 1)

OpenStudy (welshfella):

- 7x - 3 is placed in another pair of brackets

OpenStudy (welshfella):

* 7x - 3

OpenStudy (welshfella):

finally we are left with (7x - 3)(x + 2) ------------ (7x - 3)(x - 1)

OpenStudy (welshfella):

Now the 7x-3 is in top and bottom parts so it cancels out so we are left with (x + 2) / ( x - 1)

OpenStudy (welshfella):

which is choice b

OpenStudy (prince1342):

Hm. I started to understand more towards the end. I think because of the long process it confuses me. Can you help with another? It'll help me a lot, ill post it seperately of course

OpenStudy (welshfella):

I made a mistake in calculating the second part but we got the right answer in the end Sorry I have to go now for 30 minutes. be back later.

OpenStudy (prince1342):

Oh okay!

OpenStudy (prince1342):

thanks for the help

OpenStudy (welshfella):

Yeah it is a long process and takes some practice to perfect.

OpenStudy (welshfella):

yw